# RECAP: Reproducing Copyrighted Data from LLMs Training with an Agentic Pipeline

### v1

@@ 5.3 Model Size
It is well documented by prior studies that larger models tend to memorize more training data (Morris et al., 2025). Our findings are consistent with this pattern: as shown in Figure 5, all extraction methods improve as model size increases, with GPT-4.1 clearly outperforming GPT-4.1 Nano. Furthermore, RECAP consistently achieves the highest ROUGE-L scores at every model size, highlighting its effectiveness beyond existing baselines.
@@ 6 Conclusion
We tested RECAP on our new benchmark, EchoTrace, which includes research papers and books from different categories, such as public domain and copyrighted texts. While we acknowledge RECAP to be computationally intensive (Appendix T), across multiple model families, RECAP consistently outperforms all other methods; as an illustration, it extracted $\approx$3,000 passages from the first "Harry Potter" book with Claude-3.7, compared to the 75 passages identified by the best baseline.
@@
We will not release any copyrighted passages uncovered during this project. However, we plan to make publicly available all the public domain books used in EchoTrace, including the segmented passages, the event summaries, and the model outputs. We believe, therefore, our work to be fully aligned with the current legal and ethical standards for scientific research (Rosati, 2018).
@@ Appendix A Limitations
Another consideration is the prompt-intensiveness of RECAP. The iterative feedback loop inevitably increases the number of API calls compared to single-iteration approaches. Although we introduce measures such as hybrid memorization score filtering to mitigate unnecessary iterations (Section 5.7), the pipeline can still be costly, especially at scale. To put this in perspective, extracting DeepSeek-V3 memorized content from Harry Potter and the Philosopher’s Stone costs about \$2 with RECAP, compared to \$0.87 for DSP and \$0.07 for prefix-probing (Appendix T). While this makes RECAP more resource-intensive, its cost remains modest in absolute terms and still practical in smaller high-stakes settings, such as an author investigating whether their work was memorized by a model.
@@ Appendix B Section Summary Agent Details
Table 4: The Section Summary Agent, by leveraging OpenAI Structured Outputs feature, produces JSON-formatted results. Here we present an example for the book: A Christmas Carol book.
@@ Appendix E Verbatim Verifier Details
Table 9: Gemini-2.5-Flash Misclassifications for the different models: The Hobbit book (180 events/model).
@@
Despite this issue, the frequency of misclassifications remains very low. Specifically, we observed only six misclassifications across all the tested models. Given this low error rate, we believe that Gemini is a justified choice as the verbatim verifier model for this task.
@@ Appendix F Jailbreaker Details
The Jailbreaker prompt is our adaptation of the Narrative Tool Injection method, as described on the Injectprompt333www.injectprompt.com/p/claude-37-sonnet-jailbreak-narrative-tool-injection website. As before, we use the contextual information already provided to the model in the default extraction prompt, but we frame the problem differently. The aim is to make the model believe it can execute a trusted function. By instructing the model to simulate the output of this function, we are often able to overcome its typical refusals. The prompt is presented in Table 10.
@@ G.1 Feedback Agent - Real Example
Below, in Tables 12, 13, 14 and 15 we present a concrete example of the feedback process in action, using an event from Around the World in Eighty Days by Jules Verne. After just a single round of feedback, we observe a clear improvement in the model’s output, demonstrating the effectiveness of the approach.
@@ Appendix I Implementation
Our evaluation employs multiple models, including black-box ones such as Claude-3.7 (Anthropic, 2025) and GPT-4.1 (OpenAI, 2025), as well as white-box models like DeepSeek-V3 (DeepSeek-AI, 2025) and Qwen-3 (Yang et al., 2025). The latter is run locally on a computing cluster equipped with four NVIDIA A100 80GB GPUs and using vLLM [^4] https://docs.vllm.ai/. All other models are accessed via their APIs.
Generation hyperparameters are also selected to match the requirements of each task. For section-level summaries and event segmentations, we set the temperature to 1.0 to encourage stylistic diversity in the outputs. In contrast, for extraction attempts or feedback-guided generations, we use a temperature of 0.0 to ensure more completions.
@@
Figure 9: Initial refusal rates for Dynamic Soft Prompting (DSP) and the proportion of these refusals successfully bypassed with our Jailbreaking module. While baseline refusal levels vary widely across models, our jailbreak strategy consistently overcomes the vast majority of blocked cases, enabling extraction even under strong alignment safeguards.
@@
The impact of jailbreaking also remains relevant for these smaller models, though its contribution is somewhat diminished compared to the larger models. For example, enabling jailbreaking improves the average ROUGE-L from 0.242 to 0.265 on the copyrighted content (a 10% gain), where in Table 2 the relative improvement is nearly 42%. This suggests that smaller models like GPT-4.1 Nano or Qwen-3 32B are less aggressively aligned, and consequently refusals are less frequent.
@@
While our experiments employ GPT-4.1 as the default feedback model, we also evaluated whether different options could influence the overall extraction performance. However, the set of refined events varies by model, so we normalize each one’s ROUGE-L score by scaling it to the number of events refined by the model with the highest coverage (e.g., if one refines 1200 events and another 1000, the latter’s ROUGE-L is multiplied by $\frac{1000}{1200}$).
@@
Figure 11 provides a complementary perspective on the relationship between a book’s commercial success and the likelihood of its memorization by LLMs, this time focusing on smaller models. In contrast to the pronounced positive correlation observed for larger models (as seen in Figure 5), here the association between copies sold and ROUGE-L scores appears considerably more subtle. This aligns with our earlier findings in the model size analysis (Section 5.3), where we show that smaller models tend to exhibit a much more limited capacity for memorization.
@@
In Section 5.5 we focused on DeepSeek-V3 to illustrate how iterative refinement impacts extraction. Here we report the corresponding results for the other main models. Figure 12 shows that the qualitative pattern is highly similar.
@@
In a first place we see that GPT-4.1 and Gemini-2.5-Pro exhibit strong improvements after the first feedback round. However, additional iterations lead to negligible increases in ROUGE-L. Claude-3.7, on the other hand, demonstrates a more sustained improvement curve. While the first iteration still provides the majority of gains, the second round still delivers measurable additional gains, suggesting that Claude can leverage feedback corrections more effectively than the other models.
@@
Appendix Q Looking for Further Evidence of Possible Contamination in RECAP
@@
Appendix R Matching with no Tolerance for Verbatim Mismatches
As described in Section 4.1, our main analysis considered a passage as memorized if all 40 tokens matched with the gold reference ones, allowing only up to five token mismatches ($\approx 12.5\%$). This small tolerance was introduced to account for minor formatting differences such as punctuation, hyphenation, or line breaks, which do not affect whether the passage is truly memorized. Here we repeat the analysis with different mismatch thresholds, including the strict zero-tolerance case, to test the robustness of RECAP’s extraction ability.
@@
Appendix S Reducing the Number of Feedback Iterations in RECAP
S.1 Remove Verbatim Verifier and Use of Jailbreaking by Default
@@
To assess whether this approach yields practical benefits in our context, we conducted one experiment for which we present the results in Figure 17.
@@
The results indicate that the default RECAP pipeline, which includes the Verbatim Verifier, outperforms the approach that applies the Jailbreaker prompt by default. This is likely because the Jailbreaker prompt (Appendix F) is more complex, requiring the model to simulate a function call with multiple parameters, which can distract from the core extraction task and lead to lower accuracy. These findings suggest that the Jailbreaker prompt should be reserved for cases where the standard, simpler prompt fails, rather than used universally.
S.2 Hybrid Memorization Score - Choice of Thresholds
@@
Appendix T Prompt, Token, and Cost Analysis - Baselines and RECAP

### v3

@@ 5.3 Model Size
It is well documented by prior studies that larger models tend to memorize more training data (meta_memorization_work). Our findings are consistent with this pattern: as shown in Figure 5, all extraction methods improve as model size increases, with GPT-4.1 clearly outperforming GPT-4.1 Nano. Furthermore, RECAP consistently achieves the highest ROUGE-L scores at every model size, highlighting its effectiveness beyond existing baselines.
@@ 6 Conclusion
We tested RECAP on our new benchmark, EchoTrace, which includes research papers and books from different categories, such as public domain and copyrighted texts. While we acknowledge RECAP to be computationally intensive (Appendix V), across multiple model families, RECAP consistently outperforms all other methods; as an illustration, it extracted $\approx$3,000 passages from the first "Harry Potter" book with Claude-3.7, compared to the 75 passages identified by the best baseline.
@@
We will not release any copyrighted passages uncovered during this project. However, we plan to make publicly available all the public domain books used in EchoTrace, including the segmented passages, the event summaries, and the model outputs. We believe, therefore, our work to be fully aligned with the current legal and ethical standards for scientific research (tdm_eu).
@@ Appendix A Limitations
Another consideration is the prompt-intensiveness of RECAP. The iterative feedback loop inevitably increases the number of API calls compared to single-iteration approaches. Although we introduce measures such as hybrid memorization score filtering to mitigate unnecessary iterations (Section 5.7), the pipeline can still be costly, especially at scale. To put this in perspective, extracting DeepSeek-V3 memorized content from Harry Potter and the Philosopher’s Stone costs about \$2 with RECAP, compared to \$0.87 for DSP and \$0.07 for prefix-probing (Appendix V). While this makes RECAP more resource-intensive, its cost remains modest in absolute terms and still practical in smaller high-stakes settings, such as an author investigating whether their work was memorized by a model.
@@ Appendix B Section Summary Agent Details
Table 5: The Section Summary Agent, by leveraging the OpenAI Structured Outputs feature, produces JSON-formatted results. Here we present an example for the book: A Christmas Carol book.
@@ Appendix E Verbatim Verifier Details
Table 10: Gemini-2.5-Flash Misclassifications on different models: The Hobbit book (180 events/model).
@@ Appendix F Jailbreaker Details
The Jailbreaker prompt is our adaptation of the Narrative Tool Injection method, as described on the Injectprompt444www.injectprompt.com/p/claude-37-sonnet-jailbreak-narrative-tool-injection website. As before, we use the contextual information already provided to the model in the default extraction prompt, but we frame the problem differently. The aim is to make the model believe it can execute a trusted function. By instructing the model to simulate the output of this function, we are often able to overcome its typical refusals. The prompt is presented in Table 11.
@@ G.1 Feedback Agent - Real Example
Below, in Tables 13, 14, 15 and 16 we present a concrete example of the feedback process in action, using an event from Around the World in Eighty Days by Jules Verne. After just a single round of feedback, we observe a clear improvement in the model’s output, demonstrating the effectiveness of the approach.
@@ Appendix I Implementation
Our evaluation employs multiple models, including black-box ones such as Claude-3.7 (Claude-3.7) and GPT-4.1 (GPT-4.1), as well as white-box models like DeepSeek-V3 (deepseek-v3) and Qwen-3 (qwen3). The latter is run locally on a computing cluster equipped with four NVIDIA A100 80GB GPUs and using vLLM [^5] https://docs.vllm.ai/. All other models are accessed via their APIs.
Generation hyperparameters are also selected to match the requirements of each task. For section-level summaries and event segmentations, we set the temperature to 1.0 to encourage stylistic diversity in the outputs. In contrast, for extraction attempts or feedback-guided generations, we use temperature=0.
@@
Figure 9: Initial refusal rates for Dynamic Soft Prompting (DSP) and the proportion of these refusals successfully bypassed with our Jailbreaking module. While baseline refusal levels vary widely across models, our jailbreak strategy consistently overcomes the vast majority of blocked cases.
@@
The impact of jailbreaking remains relevant for smaller models, though its effect is weaker than for larger ones. For example, enabling jailbreaking increases the average ROUGE-L on copyrighted content from 0.242 to 0.265 (a 10% gain), while Table 3 shows a nearly 42% improvement for larger models. This suggests that smaller models are less aggressively aligned, resulting in fewer refusals.
@@
While our experiments use GPT-4.1 as the default feedback model, we also evaluate alternative models to assess their effect on extraction performance. Because models refine different numbers of events, we normalize ROUGE-L scores by scaling them relative to the model with the highest event coverage (e.g., if one model refines 1200 events and another 1000, the latter’s ROUGE-L is multiplied by $\frac{1000}{1200}$).
@@
Figure 11 provides a complementary view of the relationship between a book’s commercial success and its likelihood of memorization by LLMs, focusing on smaller models. Unlike the strong positive correlation observed for larger models (Figure 5), the association between copies sold and ROUGE-L scores here is much weaker. This is consistent with our earlier model size analysis (Section 5.3), which shows that smaller models have a more limited memorization capacity.
@@
In Section 5.5, we used DeepSeek-V3 to illustrate how iterative refinement impacts extraction. Here we report results for other models, which show a highly similar qualitative pattern (Figure 12).
@@
First, GPT-4.1 and Gemini-2.5-Pro exhibit strong improvements after the first feedback round. However, additional iterations lead to negligible increases in ROUGE-L. Claude-3.7, on the other hand, demonstrates a more sustained improvement curve. While the first iteration still provides the majority of gains, the second round still delivers measurable additional gains, suggesting that Claude can leverage feedback corrections more effectively than the other models.
@@
Appendix Q Ablation Study on Open-Weights Model with Open-Data
Figure 15: OLMo-2 32B book-wise performance on EchoTrace.
All the models tested so far are either closed-source or open-source models that do not clearly disclose the exact data used during training. While an educated guess suggests that most of them train on large-scale web scrapes that almost certainly include high-quality public-domain sources (e.g., Project Gutenberg), this remains only a hypothesis. For this reason, we repeat the EchoTrace experiments on a fully open model: OLMo-2-32B, which is not only open-weights but also provides transparent documentation of its training data sources olmo2.
Dolma (OLMo’s primary training corpus) contains over 12TB of text, making exact membership checking for each passage computationally infeasible dolma. Nevertheless, two properties of the dataset allow us to reason cleanly about the expected exposure of each book category in EchoTrace: (i) Dolma explicitly includes approximately 560K Project Gutenberg books, which guarantees that all public-domain titles in our benchmark are present in the training data; and (ii) in its public filing to the U.S. Copyright Office’s 2023 Notice of Inquiry on Generative AI, the Allen Institute for AI states that “much or most of our training data consists of copyrighted works" ai2_copyright_report. This suggests that several of our copyrighted bestsellers, widely redistributed online, may plausibly appear in OLMo’s training data as well. Finally, our non-training books (released in mid-2025), therefore after OLMo-2’s training cutoff, ensure they are absent.
Figure 15 shows that OLMo-2-32B displays a similar pattern as observed in Appendixes O and P.
  • (i) Strong extractability for public-domain books that are known to be in the training set,
  • (ii) “Training-like” extraction behavior for several copyrighted bestsellers.
  • (iii) No valid extractions for guaranteed unseen books.
This alignment with our previous experiments provides an additional validation for our closed-source model findings.
Appendix R Ablation Study on Low Popularity Books
Figure 16: GPT-4.1 performance on 5 low popularity books selected from Project Gutenberg.
EchoTrace focuses primarily on widely known works, which is an intentional design choice, as such texts are more likely to have been included repeatedly in the LLMs training data and therefore provide clearer memorization signals. However, this design naturally raises a further question: How does RECAP behave on texts that are substantially less likely to appear as often during training? In Section 5.4 we observed a preliminary correlation between book popularity and extractability, but this analysis did not address the extreme end of the distribution: public-domain works that are likely in the training set but receive almost no online circulation.
To investigate this regime, we conduct an additional ablation study on a set of five very low-popularity public-domain books sourced from Project Gutenberg. For context, at the time of writing, Frankenstein exceeds 250k downloads over the past month. By contrast, each of the books chosen for this ablation records fewer than 500 monthly downloads, placing them in the long tail of the Gutenberg catalog and making them far less likely to have been frequently encountered during pretraining.
We evaluate these books using GPT-4.1, which has shown to be capable of memorizing substantial amounts of training data (Appendix O), and apply the full RECAP pipeline without modifications.
As Figure 16 illustrates, the number of extracted passages across all five low-popularity books is markedly lower than what we observe for popular public-domain texts (Appendix O). This is exactly the behavior we would expect if these books were either never included in training or appeared only minimally in the crawled data.
These results reinforce a point we already emphasized in Appendix A: A strong RECAP signal offers evidence of memorization, but a weak signal should not be interpreted as proof of non-membership.
Appendix S Looking for Further Evidence of Possible Contamination in RECAP
@@
Appendix T Matching with no Tolerance for Verbatim Mismatches
As described in Section 4.1, our main analysis considers a passage memorized if all 40 tokens match the gold reference, allowing up to five mismatches to account for minor formatting differences (e.g., punctuation). We repeat the analysis with different mismatch thresholds to test the robustness of RECAP.
@@
Appendix U Reducing the Number of Feedback Iterations in RECAP
U.1 Remove Verbatim Verifier and Use of Jailbreaking by Default
@@
Figure 19 shows that the default RECAP pipeline, which includes the Verbatim Verifier, outperforms applying the Jailbreaker prompt by default. This likely occurs because the Jailbreaker prompt (Appendix F) is more complex, requiring the model to simulate a function call with multiple parameters, which can distract from the extraction task and reduce performance. Thus, although Appendix J shows that the Jailbreaker reliably converts most refusals into valid completions, it is most effective when applied conditionally rather than by default.
U.2 Hybrid Memorization Score - Choice of Thresholds
@@
Appendix V Prompt, Token, and Cost Analysis - Baselines and RECAP